# Frequency-Guided Posterior Sampling for Diffusion-Based Image Restoration

### v1

@@ 5 Frequency Guided Posterior Sampling
While the results from Figures 1 and 2 might appear artificial, this phenomenon frequently emerges in complex real-world image restoration problems. For example, in directional motion deblurring, the forward operator does indeed act as a high-pass filter in certain spatial directions of the image (see Appendix C). Unfortunately, for real image data that follows a complex non-Gaussian distribution, we cannot exactly compute the true conditional score as we did in Theorem 2. Despite this, we hope to derive practical insights from our theoretical analysis because the underlying data distribution in our synthetic experiments mimics the frequency distribution of natural images.
Specifically, Figure 2 reveals a crucial insight: when using a low-pass filter as the forward operator, the approximation gap remains small during the early stages of the reverse process. Diffusion models inherently generate images hierarchically, from coarse to fine details, so the early stages of the reverse process play a crucial role in establishing the fundamental structure that matches the measurements. If the approximation gap is large in this regime, as is the case when the forward operator is a high-pass filter, it is unlikely the sampling procedure will produce a good reconstruction.
We propose a simple fix to this issue by constraining the measurements to retain only low frequency components at the initial steps of the reverse process and progressively add in high frequency components. This intuitively aligns the frequency characteristics of the measurements with the iterates of the diffusion reverse process. To implement this, we convolve the original measurement $\mathbf{y}$ with a time-varying low-pass filter $\mathbf{k}_{t}$. Denoting $\phi_{t}(\mathbf{y})$ as this convolution, the approximation we propose for the noisy likelihood score is
@@
where $\bm{\mu}_{0\mid t}$ denotes the Tweedie estimate $\mathbb{E}[\mathbf{x}_{0}\mid\mathbf{x}_{t}]$ predicted by the diffusion model. In order to compute the above approximation, we need to compute the model likelihood for the new measurements. Since convolution of a signal $\mathbf{y}$ with a low pass filter can be represented as a circulant matrix multiplied by $\mathbf{y}$, we have that $\phi_{t}(\mathbf{y})=\mathbf{C}_{t}\mathbf{y}$ where $\mathbf{C}_{t}$ is a time-dependent circulant matrix representing $\mathbf{k}_{t}$. Thus, using properties of the multivariate normal distribution, we have that the updated model likelihood is
@@
Dataset-informed choice of $\mathbf{k}_{t}$. An important choice for our method is choosing the sequence $\{\mathbf{k}_{t}\}_{t=1}^{T}$. Each $\mathbf{k}_{t}$ is a low-pass filter that should allow frequencies up till a threshold $\tau_{t}$ and then strongly attenuate any frequencies above $\tau_{t}$. In order to encode that the initial steps of the reverse process should only use low frequency components of $\mathbf{y}$, we require that $\tau_{T}$ should start off small and increase as $t$ decreases. We denote the sequence of $\tau_{t}$ as our frequency curriculum. As datasets can have varying frequency characteristics, we set our frequency curriculum in a data-dependent way, as explained in Section 6.
Implementation Details. We efficiently compute $\phi_{t}(\cdot)$ in the Fourier domain. Specifically, $\mathbf{C}_{t}$ can be diagonalized as $\mathbf{C}_{t}=\mathbf{F^{-1}MF}$, where $\mathbf{F}$ denotes the Discrete Fourier Transform matrix [37]. Using this property along with the fact that $\mathbf{F}$ is an orthogonal matrix, we have that $p(\phi_{t}(\mathbf{y})\mid\bm{\mu}_{0\mid t})=p(\mathbf{F}\phi_{t}(\mathbf{y})\mid\bm{\mu}_{0\mid t})=p(\mathbf{MFy}\mid\bm{\mu}_{0\mid t})$. Thus, we compute $\phi_{t}(\mathbf{y})$ and $\phi_{t}(\mathcal{A}(\bm{\mu}_{0\mid t}))$ by first taking the Fast Fourier Transform of the input and applying a mask in the frequency domain, which adds minimal computational overhead (see Appendix D). For simplicity, the mask for filter $\phi_{t}(\cdot)$ is a binary mask that zeroes out the frequency components larger than $\tau_{t}$.
@@
1:$\mathbf{y}$, $\mathcal{A}$, $\{\phi_{t}(\cdot),\mathbf{S}_{t},\alpha_{t},\tilde{\sigma}_{t}\}_{t=1}^{T}$
2:$\mathbf{x}_{T}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$
3:for $t\leftarrow T$ to $1$ do
4:     $\bm{\hat{s}}\leftarrow s_{\theta}(\mathbf{x}_{t},t)$
@@
6:     $\bm{\mu}_{0\mid t}\leftarrow\frac{1}{\sqrt{\bar{\alpha}_{t}}}(\mathbf{x}_{t}+(1-\bar{\alpha}_{t})\bm{\hat{s}})$
@@
8:     $\mathbf{x}_{t-1}^{\prime}\leftarrow\frac{\sqrt{\alpha_{t}}(1-\bar{\alpha}_{t-1})}{1-\bar{\alpha}_{t}}\mathbf{x}_{t}+\frac{\sqrt{\bar{\alpha}_{t-1}}(1-\alpha_{t})}{1-\bar{\alpha}_{t}}\bm{\mu}_{0\mid t}+\tilde{\sigma}_{t}\mathbf{z}$
9:     $\mathbf{x}_{t-1}\leftarrow\mathbf{x}_{t-1}^{\prime}-\mathbf{S}_{t}\nabla_{\mathbf{x}_{t}}\left\|\phi_{t}(\mathbf{y})-\phi_{t}(\mathcal{A}(\bm{\mu}_{0\mid t}))\right\|_{2}^{2}$
@@ 6 Experiments
To empirically validate the performance of FGPS, we begin by investigating the approximation gap of our method on synthetic data as in Figure 2. Then, we evaluate our method on several real-world datasets and image restoration tasks, where we demonstrate compelling performance of our method as compared to state-of-the-art baselines.
6.1 Synthetic Data: FGPS Approximation Gap
Figure 3: We use the same setup as Figure 2. Our method has a significantly smaller approximation gap than DPS when the forward operator is a high-pass filter. Note the $y$-axis is in log-scale.
From Theorem 2, we know that under a multivariate normal assumption on the data, the posterior mean $\bm{\mu}_{0\mid t}$ is $\bm{\Gamma}_{t}\mathbf{x}_{t}$. Under the same assumptions as Theorem 2, for linear inverse problems where $\mathcal{A}(\mathbf{x}_{0})=\mathbf{A}\mathbf{x}_{0}$ we analytically compute our approximation to the noisy likelihood score given in Equation 15 as
Theorem 2 gives an exact expression for the true noisy likelihood score, so the approximation gap of our method can also be computed analytically. In Figure 3, we plot the approximation gap of our method (with frequency curriculum given in Appendix D) as compared to DPS. We find that in the early steps of the reverse process, our method has an approximation gap that is several orders of magnitude smaller than DPS when the forward operator is a high-pass filter. As the early stages of the reverse process are when the generated image develops coarse features, we argue it is crucial that the approximation gap is small in these phases.
@@ 6.1 Synthetic Data: FGPS Approximation Gap
Table 1: Quantitative comparison of different methods on 1000 images from the FFHQ and Imagenet validation datasets. Bolded entries indicate best performance and underlined entries indicate second best performance.
@@ 6.2 Real-World Data
Tasks. We consider four challenging image restoration tasks for our experimental evaluation. The first three are linear inverse problems: Gaussian deblurring, motion deblurring, and deconvolution with a high-pass filter. The last is a nonlinear inverse problem: image dehazing. The details of the four forward operators we consider can be found in Appendix D. All measurements have Gaussian measurement noise added with standard deviation $0.05$.
@@
Baselines and Metrics. We compare against several competitive unsupervised methods. The first two baselines are popular unsupervised diffusion-based solvers: Diffusion Posterior Sampling (DPS) [9], a method that approximates the conditional score with a Dirac delta around the posterior mean, and Score-SDE/ILVR [8, 35], a method that uses a sequence of noisy measurements computed using the diffusion model variance schedule. Then, we compare to two state-of-the-art methods focused on mitigating the approximation gap problem: Manifold Constrained Gradient (MCG) [10], which after each denoising timestep performs a projection onto an estimate of the noisy data manifold, and Diffusion with Spherical Guidance (DSG) [41], which restricts the guidance step using a spherical Gaussian constraint. For image dehazing, we further compare to two problem-specific baselines: DoubleDIP [18], an unsupervised method which uses a dark channel prior to estimate the dehazed image, and AOD-Net [26], a supervised method trained to dehaze images from pairs of clean-hazy images. For fairness, we use the same pretrained score function for all diffusion-based methods. We evaluate our method on two perceptual metrics, FID and LPIPS distance, and two distortion-based metrics, PSNR and SSIM.
Frequency Curriculum. For both FFHQ and ImageNet across all tasks, in our method we set $\tau_{1}$ and $\tau_{T}$ to be $\frac{75*100}{256}\%$ and $\frac{10*100}{256}\%$ of the overall frequency range respectively. We set different frequency curricula for the two datasets as described in Section 5. Specifically, our experiments (shown in Appendix F) reveal that on deblurring tasks on FFHQ, our method performs better with an exponentially increasing schedule between $\tau_{T}$ and $\tau_{1}$, while on ImageNet, our method achieves better results with a linear schedule.
@@
Step Size Schedule. Similar to DPS, we replace the theoretical quantity $\mathbf{S}_{t}$ for the conditional sampling step with $\frac{\kappa}{\left\|\phi_{t}(\mathbf{y})-\phi(\mathcal{A}(\bm{\mu}_{0\mid t}))\right\|_{2}}\mathbf{I}$ for a hyperparameter $\kappa$, which works well in practice. Using the intuition from Figure 3 that the approximation gap is small in the initial steps of the reverse process, we set $\kappa$ large initially and decay it following a cosine curve.

### v2

@@ 4 Frequency Guided Posterior Sampling
As illustrated in Section 2.2, the main challenge in developing unsupervised inverse problem solvers is estimating the noisy likelihood score $\nabla_{\mathbf{x}_{t}}\log p(\mathbf{y}\mid\mathbf{x}_{t})$, which depends on the intractable denoising distribution $p(\mathbf{x}_{0}\mid\mathbf{x}_{t})$. It is common to substitute this denoising distribution with an approximation that depends only on the posterior mean $\bm{\mu}_{0\mid t}=\mathbb{E}[\mathbf{x}_{0}\mid\mathbf{x}_{t}]$, which is tractable using Tweedie’s formula. For example, Diffusion Posterior Sampling (DPS) approximates the noisy likelihood score with $\nabla_{\mathbf{x}_{t}}\log p(\mathbf{y}\mid\bm{\mu}_{0\mid t})$ [9]. However, these methods can incur a significant approximation error by using simple distributions centered around the posterior mean instead of the true distribution.
The key challenge is ensuring that $\mathbf{x}_{t}$ stays on the data manifold learned by the diffusion model even after it is updated using an approximation of the gradient $\nabla_{\mathbf{x}_{t}}\log p(\mathbf{y}\mid\mathbf{x}_{t})$. Prior approaches address this by assuming that the data manifold is linear or that $\bm{\mu}_{0\mid t}$ remains on the clean data manifold [10, 42]. However, these assumptions are often violated in practice. Our main insight is to use the observation that diffusion models generate images hierarchically. Specifically, recent work has connected diffusion models and inverse heat dissipation, concluding that diffusion models naturally generate images from low frequencies (corresponding to coarse features) to high frequencies (corresponding to fine-grained features) [30]  [^1] This effect is because the forward process noise schedule removes high frequency components proportional to the variance of the noise.. If there is a mismatch in frequency content between $\bm{\mu}_{0\mid t}$ and the measurement $\mathbf{y}$, updating $\mathbf{x}_{t}$ by following an approximation of $\nabla_{\mathbf{x}_{t}}\log p(\mathbf{y}\mid\mathbf{x}_{t})$ can move $\mathbf{x}_{t}$ off the learned data manifold. This problem is exacerbated during the early stages of the reverse process since the early stages play a crucial role in ensuring the generated output matches the fundamental low-frequency structure of the measurements.
We propose a simple fix to this issue by retaining only low frequency components of the measurements at the initial steps of the reverse process and progressively adding high frequency components in later steps. To implement this, we convolve the original measurement $\mathbf{y}$ with a time-varying low-pass filter $\mathbf{k}_{t}$. Denoting $\phi_{t}(\mathbf{y})$ as this convolution, the approximation we propose for the noisy likelihood score is
@@
where $\bm{\mu}_{0\mid t}$ denotes the Tweedie estimate $\mathbb{E}[\mathbf{x}_{0}\mid\mathbf{x}_{t}]$ predicted by the diffusion model. In order to compute the above approximation, we need to compute the model likelihood for the new measurements. Since convolution of a signal $\mathbf{y}$ with a low pass filter can be represented as a circulant matrix multiplied by $\mathbf{y}$, we have that $\phi_{t}(\mathbf{y})=\mathbf{C}_{t}\mathbf{y}$ where $\mathbf{C}_{t}$ is a time-dependent circulant matrix representing $\mathbf{k}_{t}$. This circulant matrix has its first row as the flattened $\mathbf{k}_{t}$ and all other rows as cyclic shifts. Thus, using properties of the multivariate normal distribution, we have that the updated model likelihood is
@@
Dataset-informed choice of $\mathbf{k}_{t}$. An important choice for our method is choosing the sequence $\{\mathbf{k}_{t}\}_{t=1}^{T}$. Each $\mathbf{k}_{t}$ is a low-pass filter that should allow frequencies up till a threshold $\tau_{t}$ and then strongly attenuate any frequencies above $\tau_{t}$. In order to encode that the initial steps of the reverse process should only use low frequency components of $\mathbf{y}$, we require that $\tau_{T}$ should start off small and increase as $t$ decreases. We denote the sequence of $\tau_{t}$ as our frequency curriculum. As datasets can have varying frequency characteristics, our method gives flexibility to set the frequency curriculum in a data-dependent way, as explained in Section˜6.
Efficient Implementation. We efficiently compute $\phi_{t}(\cdot)$ in the Fourier domain. Specifically, $\mathbf{C}_{t}$ can be diagonalized as $\mathbf{C}_{t}=\mathbf{F^{-1}MF}$, where $\mathbf{F}$ denotes the Discrete Fourier Transform matrix [39]. Using this property along with the fact that $\mathbf{F}$ is an orthogonal matrix with determinant $1$, we have that $p(\phi_{t}(\mathbf{y})\mid\bm{\mu}_{0\mid t})=p(\mathbf{F}\phi_{t}(\mathbf{y})\mid\bm{\mu}_{0\mid t})=p(\mathbf{MFy}\mid\bm{\mu}_{0\mid t})$. Thus, we bypass storing the large circulant matrix and compute $\phi_{t}(\mathbf{y})$ and $\phi_{t}(\mathcal{A}(\bm{\mu}_{0\mid t}))$ by first taking the Fast Fourier Transform of the input and applying a binary mask in the frequency domain, which adds minimal computational overhead (see Appendix E).
@@
[…]
3:for $t\leftarrow T$ to 11 do
4:  $\bm{\hat{s}}\leftarrow s_{\theta}(\mathbf{x}_{t},t)$ // Predict $\nabla_{\mathbf{x}_{t}}\log p(\mathbf{x}_{t})$
@@
[…]
@@
8:  $\mathbf{x}_{t-1}^{\prime}\leftarrow\frac{\sqrt{\alpha_{t}}(1-\bar{\alpha}_{t-1})}{1-\bar{\alpha}_{t}}\mathbf{x}_{t}+\frac{\sqrt{\bar{\alpha}_{t-1}}(1-\alpha_{t})}{1-\bar{\alpha}_{t}}\bm{\mu}_{0\mid t}+\tilde{\sigma}_{t}\mathbf{z}$ // DDIM Sampling
9:  $\mathbf{x}_{t-1}\leftarrow\mathbf{x}_{t-1}^{\prime}-\mathbf{S}_{t}\nabla_{\mathbf{x}_{t}}\left\|\phi_{t}(\mathbf{y})-\phi_{t}(\mathcal{A}(\bm{\mu}_{0\mid t}))\right\|_{2}^{2}$ // FGPS Approximation of $\nabla_{\mathbf{x}_{t}}\log p(\mathbf{y}\mid\mathbf{x}_{t})$
@@
5 Theoretical Analysis
In this section, we utilize distributional assumptions to provide a precise theoretical analysis of the approximation errors induced by traditional posterior sampling methods as well as our proposed method.
5.1 Mind the (Approximation) Gap
Since we rely on a spectral characterization of the reverse diffusion process, our analysis requires two key properties.
  1. 1.
  2. 2.
For the first property, we look at the power spectral density of the data, as defined in Equation˜8, which tells us how much the signal’s power is distributed across different frequencies. For the second property, it is natural to suppose the clean data $\mathbf{x}_{0}$ comes from a multivariate Gaussian distribution. Although these two properties seem conflicting, the following lemma establishes an equivalence between the data power spectral density and the data covariance.
Lemma 1.
Let $f_{k}=\frac{k}{n}$ for $k=0,\dots,n-1$ denote the DFT sample frequencies for a signal of length $n$. Then, there exists a covariance matrix $\bm{\Sigma}_{f}$ such that for $\mathbf{x}\sim\mathcal{N}(\mathbf{0},\bm{\Sigma}_{f})$, the following two properties hold. First, the signal $\mathbf{x}$ follows a power law in the frequency domain with parameters $c,\beta>0$, i.e., it has a power spectral density $S(f_{k})=c|f_{k}|^{-\beta}$ for the non-zero DFT sample frequencies $f_{k}$. Second, the eigenvalues of $\bm{\Sigma}_{f}$ are precisely $c|f_{k}|^{-\beta}$.
This is a special case of a classical result known as the Wiener-Khinchin Theorem [7]. Note that this does not imply that all random vectors that follow a power law in the frequency domain are Gaussian. For example, natural image data tends to follow a radially averaged power law in the frequency domain, but is far from Gaussian [34]. Nevertheless, this lemma gives a distributional assumption that mimics the spectral properties of natural image data, which we will show already elucidates failure cases of existing methods. Next, we use these assumptions to exactly quantify the approximation gap of our method.
Theorem 2.
Suppose $\mathbf{x}_{0}$ is drawn from $\mathcal{N}(\mathbf{0},\bm{\Sigma})$ and we are given linear measurements $\mathbf{y}=\mathcal{A}(\mathbf{x}_{0})+\mathbf{z}$, where $\mathcal{A}(\mathbf{x}_{0})=\mathbf{A}\mathbf{x}_{0}$ and $\mathbf{z}\sim\mathcal{N}(0,\sigma_{y}^{2}I)$. Suppose that the intermediate value $\mathbf{x}_{t}$ of the continuous-time reverse diffusion from Equation˜5 can be written as $\mathbf{x}_{t}=\sqrt{\bar{\alpha}_{t}}\mathbf{x}_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon$ where $\epsilon\sim\mathcal{N}(\mathbf{0},\mathbf{I})$. Then, we have that the true noisy likelihood score $\nabla_{\mathbf{x}_{t}}\log p(\mathbf{y}\mid\mathbf{x}_{t})$ is
$(\mathbf{A}\bm{\Gamma}_{t})^{T}\!(\mathbf{A}\bm{\Sigma}_{0\mid t}\mathbf{A}^{T}\!+\sigma_{y}^{2}\bm{I})^{-1}(\mathbf{y}-\mathbf{A}\bm{\mu}_{0\mid t}),\!\!$ (13)
where $\bm{\Gamma}_{t}=\sqrt{\bar{\alpha}_{t}}\bm{\Sigma}(\bar{\alpha}_{t}\bm{\Sigma}+(1-\bar{\alpha}_{t})\bm{I})^{-1}$, $\bm{\mu}_{0\mid t}=\mathbb{E}[\mathbf{x}_{0}\mid\mathbf{x}_{t}]=\bm{\Gamma}_{t}\mathbf{x}_{t}$ and $\bm{\Sigma}_{0\mid t}=\mathrm{Cov}[\mathbf{x}_{0}\mid\mathbf{x}_{t}]=\bm{\Sigma}-\sqrt{\bar{\alpha}_{t}}\bm{\Gamma}_{t}\bm{\Sigma}$. Moreover, the FGPS approximation $\nabla_{\mathbf{x}_{t}}\log p(\mathbf{C}_{t}\mathbf{y}\mid\bm{\mu}_{0\mid t})$ can be analytically calculated as
$\left(\mathbf{C}_{t}\mathbf{A}\bm{\Gamma}_{t}\right)^{T}(\sigma_{y}^{2}\mathbf{C}_{t}\mathbf{C}_{t}^{T})^{-1}(\mathbf{C}_{t}\mathbf{y}-\mathbf{C}_{t}\mathbf{A}\bm{\mu}_{0\mid t}).$ (14)
The proof is in Appendix B. The difference of Equation˜13 and Equation˜14 gives a precise expression for the approximation gap of our method. The $\bar{\alpha}_{t}$ corresponds to the variance schedule of standard diffusion models such as DDPM [20]. Our assumption that $\mathbf{x}_{t}=\sqrt{\bar{\alpha}_{t}}\mathbf{x}_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon$ isolates the effects of the approximation gap as the generated sample is exactly $\mathbf{x}_{0}$. Our framework also allows us to analytically study the approximation gap for other inverse problem solvers. Below, we give one example for the popular Diffusion Posterior Sampling (DPS) solver.
Corollary 3.
Under the same assumptions and notation as in Theorem˜2, we have that the DPS approximation $\nabla_{\mathbf{x}_{t}}\log p(\mathbf{y}\mid\bm{\mu}_{0\mid t})$ can be analytically calculated as
$\sigma_{y}^{-2}(\mathbf{A}\bm{\Gamma}_{t})^{T}(\mathbf{y}-\mathbf{A}\bm{\mu}_{0\mid t}).\!\!$ (15)
To the best of our knowledge, this is the first time any distributional assumptions have been used to characterize the exact approximation gap of any inverse problem solver. While [9] and [42] provide upper and lower bounds on this gap, these bounds do not exploit the structure of the data distribution and as such, can be loose. We formally compare the approximation gap of FGPS and DPS in Appendix C.
5.2 Approximation Gap on Synthetic Data
Figure 1: We show the exact approximation gap for FGPS and DPS across diffusion timesteps when the data follows a power law in the frequency domain and the forward operator is a a high-pass filter (Dirac kernel minus a Gaussian kernel) of varying width. The width denotes the $\sigma$ of the Gaussian kernel. Note the $y$-axis is in log-scale.
In this subsection, we probe the approximation gap of our method and DPS on synthetic data, and we give examples of forward operators where FGPS results in a significantly less approximation gap than DPS. In order to construct these examples, we note that Theorem˜2 depends crucially on how the forward operator $\mathbf{A}$ interacts with the data covariance $\bm{\Sigma}$, particularly through the term $\mathbf{A}\bm{\Sigma}_{0\mid t}\mathbf{A}^{T}$. Intuitively, if the data covariance is chosen to be $\bm{\Sigma}_{f}$ from Lemma 1, the frequency characteristics of the data and the forward operator would be misaligned when the forward operator is a high-pass filter, so the measurement contains primarily high-frequency components. Some natural imaging systems fall into this regime, e.g., phase contrast microscopy where intensity measurements are high-pass filtered versions of refractive index.
To verify the above intuition, we begin by simulating 1-D signals $\mathbf{x}_{0}\sim\mathcal{N}(\mathbf{0},\bm{\Sigma}_{f})$ that follow a power law in the frequency domain, which uses the construction of $\bm{\Sigma}_{f}$ given in Lemma 1. Then, the corresponding measurements for these signals are $\mathbf{y}=\mathbf{A}\mathbf{x}_{0}+\mathbf{z}$ with $\mathbf{z}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ , where $\mathbf{A}$ is the convolution matrix corresponding to convolution with a Dirac kernel minus a Gaussian kernel of standard deviation $\sigma$. This kernel approximates a high-pass filter.
Figure 1 shows the approximation gap of DPS and FGPS as a function of the diffusion timesteps (where higher timesteps correspond to a lower value of $\bar{\alpha}_{t}$) for the variance schedule used in the DDPM model for $T=1000$ [20]. In accordance with our intuition, we see that the DPS approximation worsens significantly as we increase $t$, likely because higher frequency noise components in $\mathbf{x}_{t}$ are amplified by the high-pass filter. In contrast, we find our method has an approximation gap that is several orders of magnitude smaller than DPS in the early stages of the reverse process. As the early stages of the reverse process are when the generated image develops coarse features, we argue it is crucial that the approximation gap is small in these phases. Refer to Appendix E for the full experimental details.
@@ 6 Experiments
Although our theoretical results were restricted to cases where FGPS provably improved over existing methods, we now turn to empirical validation of FGPS on several real-world datasets and image restoration tasks.
@@
Table 1: Quantitative comparison of different methods on FFHQ and ImageNet validation datasets. Bolded entries indicate best performance and underlined entries indicate second best performance.
@@ 6.1 Experimental Setup
Tasks. We consider four challenging image restoration tasks for our experimental evaluation. The first three are linear inverse problems: Gaussian deblurring, motion deblurring, and deconvolution with a high-pass filter. The last is a nonlinear inverse problem: image dehazing. The details of the four forward operators we consider can be found in Appendix E. All measurements have Gaussian measurement noise added with standard deviation $0.05$.
@@
Baselines and Metrics. We compare against several competitive unsupervised methods. The first two baselines are popular unsupervised diffusion-based solvers: Diffusion Posterior Sampling (DPS) [9], a method that approximates the conditional score with a Dirac delta around the posterior mean, and Score-SDE/ILVR [8, 37], a method that uses a sequence of noisy measurements computed using the diffusion model variance schedule. Then, we compare to two state-of-the-art methods focused on mitigating the approximation gap problem: Manifold Constrained Gradient (MCG) [10] and Diffusion with Spherical Guidance (DSG) [42]. For image dehazing, we further compare to two problem-specific baselines: DoubleDIP [18] and AOD-Net [26], a supervised method trained to dehaze images from pairs of clean-hazy images. We use the same pretrained score function for all diffusion-based methods. We evaluate our method on two perceptual metrics, FID and LPIPS distance, and two distortion-based metrics, PSNR and SSIM.
Frequency Curriculum. For both FFHQ and ImageNet across all tasks, in our method we set $\tau_{1}$ and $\tau_{T}$ to be $\frac{75*100}{256}\%$ and $\frac{10*100}{256}\%$ of the overall frequency range respectively. We set different frequency curricula for the two datasets as described in Section˜4. We find the exponential schedule works well for face datasets with many high-frequency details, while the uniform schedule is better for diverse, multiclass datasets like ImageNet. We emphasize that $\mathbf{k}_{t}$ does not require extensive tuning, as a simple schedule already beats baseline methods (see Appendix G).
@@
Step Size Schedule. In practice, we take $\mathbf{S}_{t}$ to be a hyperparameter as opposed to being fixed as $(\sigma_{y}^{2}\mathbf{C}_{t}\mathbf{C}_{t}^{T})^{-1}$ in order to further control the effect of the approximation error. We adopt the choice that DPS makes where the step size is $\frac{\kappa}{\left\|\phi_{t}(\mathbf{y})-\phi(\mathcal{A}(\bm{\mu}_{0\mid t}))\right\|_{2}}\mathbf{I}$ for a hyperparameter $\kappa$. Using the intuition from Figure˜1 that the approximation gap is small in the initial steps of the reverse process, we set $\kappa$ large initially and decay it following a cosine curve.
6.2 Results
Empirical Validation of Theoretical Results. In Figure 2, we consider two forward operators on an example from the FFHQ dataset: a) a standard Gaussian blurring low-pass filter, and b) the high-pass filter we studied in our theoretical analysis. As expected, prior approaches such as Diffusion Posterior Sampling [9] and variants [10, 42] perform well at Gaussian deblurring. However, when the forward operator is a high-pass filter, remarkably these approaches all fail to produce coherent outputs, which aligns with our theoretical results about the large approximation gap of existing methods. FGPS on the other hand is still able to produce high-quality reconstructions.